# 2]Key Laboratory of System Control and Information Processing, Ministry of Education, Shanghai 200240, China

### v2

@@ 4 Numerical experiments
Pink noise Margenau-Hill 3.03233.03233.0323 5.27665.27665.2766 3.52373.52373.5237 4.33064.33064.3306
Kirkwood-Rihaczek 2.34002.34002.3400 3.74103.74103.7410 3.34313.34313.3431 4.21104.21104.2110
Born-Jordan 4.61904.61904.6190 7.07167.07167.0716 4.13394.13394.1339 5.14945.14945.1494
Page 2.04332.04332.0433 3.73703.73703.7370 3.21093.21093.2109 3.59633.59633.5963
Adaptive radial Gaussian kernel 3.47583.47583.4758 7.31207.31207.3120 3.67363.67363.6736 3.52163.52163.5216
Wiener filter 8.83148.83148.8314 10.319210.319210.3192 $\mathbf{8.0556}$ $\mathbf{7.8716}$
Adaptive CCTFD $\mathbf{18.5390}$ $\mathbf{12.8503}$ 6.19786.19786.1978 6.57996.57996.5799
Blue noise Margenau-Hill 1.58651.58651.5865 7.62287.62287.6228 3.64683.64683.6468 7.23777.23777.2377
Kirkwood-Rihaczek 0.5454-0.5454- 0.5454 6.81086.81086.8108 2.94692.94692.9469 5.42785.42785.4278
Born-Jordan 4.15574.15574.1557 8.37578.37578.3757 5.03425.03425.0342 13.017913.017913.0179
Page 0.2493-0.2493- 0.2493 6.47646.47646.4764 2.77082.77082.7708 2.32972.32972.3297
Adaptive radial Gaussian kernel 3.50463.50463.5046 7.90757.90757.9075 3.68133.68133.6813 3.20163.20163.2016
Wiener filter 8.10828.10828.1082 9.89379.89379.8937 8.68998.68998.6899 14.400014.400014.4000
@@
Red noise Margenau-Hill 2.38882.38882.3888 6.39986.39986.3998 3.63603.63603.6360 5.25625.25625.2562
Kirkwood-Rihaczek 1.77741.77741.7774 5.94565.94565.9456 3.48813.48813.4881 5.15595.15595.1559
Born-Jordan 4.01114.01114.0111 7.41677.41677.4167 4.38054.38054.3805 6.63486.63486.6348
Page 1.36061.36061.3606 5.98635.98635.9863 3.09253.09253.0925 3.86283.86283.8628
Adaptive radial Gaussian kernel 2.99442.99442.9944 6.62676.62676.6267 3.75523.75523.7552 3.40613.40613.4061
Wiener filter 11.883211.883211.8832 $\mathbf{16.7052}$ $\mathbf{13.1156}$ $\mathbf{11.9071}$
Adaptive CCTFD $\mathbf{20.4654}$ 16.149516.149516.1495 7.04967.04967.0496 7.20567.20567.2056
@@ 5 Conclusion
The convolution type of CCTFD time-frequency analysis method suitable for the adaptive filter denoising for additive noises jamming signals under the condition of low SNR was established, giving rise to the adaptive kernel function which minimizes the MSE in the WVD domain. The proposed adaptive CCTFD can automatically adjust its kernel function according to the change of signal to adapt to different signal characteristics. It turns out that its denoising effect is superior not only to some classical fixed kernel function-based CCTFDs, including the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution, and the Page distribution, but also to the adaptive radial Gaussian kernel function-based CCTFD. Under white Gaussian noise, our method outperforms the ordinary Wiener filter, while under colored noise, it performs comparably to the Wiener filter.
[…]
Acknowledgements This work was supported by the National Natural Science Foundation of China (Grant No 61901223), the Jiangsu Planned Projects for Postdoctoral Research Funds (Grant No 2021K205B), the Foundation of Key Laboratory of System Control and Information Processing, Ministry of Education (Grant No Scip20240121), and the Foundation of Key Laboratory of Computational Science and Application of Hainan Province (Grant No JSKX202401).

### v3

@@ 4 Numerical experiments
Pink noise Margenau-Hill $3.0323$ $5.2766$ $3.5237$ $4.3306$
Kirkwood-Rihaczek $2.3400$ $3.7410$ $3.3431$ $4.2110$
Born-Jordan $4.6190$ $7.0716$ $4.1339$ $5.1494$
Page $2.0433$ $3.7370$ $3.2109$ $3.5963$
Adaptive radial Gaussian kernel $3.4758$ $7.3120$ $3.6736$ $3.5216$
Wiener filter $8.8314$ $10.3192$ $\mathbf{8.0556}$ $\mathbf{7.8716}$
Adaptive CCTFD $\mathbf{18.5390}$ $\mathbf{12.8503}$ $6.1978$ $6.5799$
Blue noise Margenau-Hill $1.5865$ $7.6228$ $3.6468$ $7.2377$
Kirkwood-Rihaczek 0.5454-0.5454 $6.8108$ $2.9469$ $5.4278$
Born-Jordan $4.1557$ $8.3757$ $5.0342$ $13.0179$
Page 0.2493-0.2493 $6.4764$ $2.7708$ $2.3297$
Adaptive radial Gaussian kernel $3.5046$ $7.9075$ $3.6813$ $3.2016$
Wiener filter $8.1082$ $9.8937$ $8.6899$ $14.4000$
@@
Red noise Margenau-Hill $2.3888$ $6.3998$ $3.6360$ $5.2562$
Kirkwood-Rihaczek $1.7774$ $5.9456$ $3.4881$ $5.1559$
Born-Jordan $4.0111$ $7.4167$ $4.3805$ $6.6348$
Page $1.3606$ $5.9863$ $3.0925$ $3.8628$
Adaptive radial Gaussian kernel $2.9944$ $6.6267$ $3.7552$ $3.4061$
Wiener filter $11.8832$ $\mathbf{16.7052}$ $\mathbf{13.1156}$ $\mathbf{11.9071}$
Adaptive CCTFD $\mathbf{20.4654}$ $16.1495$ $7.0496$ $7.2056$
To further validate the practical effectiveness of the proposed adaptive CCTFD, we conduct experiments on real-world data. Specifically, the RadChar dataset is used, focusing on the RadChar-Tiny subset. A segment of the first $7500$ samples is extracted and contaminated with additive white Gaussian noise under three SNR levels: $1\mathrm{dB}$, $2\mathrm{dB}$, and $3\mathrm{dB}$. The MSE and PSNR results are summarized in Table 3. As observed, the proposed adaptive CCTFD consistently achieves the lowest MSE and highest PSNR across all SNR levels, demonstrating superior denoising performance. These results indicate that the proposed adaptive CCTFD is effective not only for simulated signals but also for real radar data, highlighting its practical applicability and robustness.
Table 3: The MSE and PSNR of seven filtering methods for the RadChar dataset under white Gaussian noise.
Noise level (SNR) 3-3 2-2 1-1
MSE Margenau-Hill $0.0639$ $0.0567$ $0.0493$
Kirkwood-Rihaczek $0.0640$ $0.0574$ $0.0509$
Born-Jordan $0.0491$ $0.0406$ $0.0319$
Page $0.1006$ $0.0965$ $0.0924$
Adaptive radial Gaussian kernel $0.1176$ $0.0983$ $0.0955$
Wiener filter 0.0169-0.0169 0.0407-0.0407 0.0672-0.0672
Adaptive CCTFD $\mathbf{-0.0317}$ $\mathbf{-0.0500}$ $\mathbf{-0.0675}$
PSNR Margenau-Hill $3.7697$ $3.8163$ $3.8651$
Kirkwood-Rihaczek $3.5877$ $3.6311$ $3.6764$
Born-Jordan $3.8719$ $3.9283$ $3.9876$
Page $3.5575$ $3.5916$ $3.6258$
Adaptive radial Gaussian kernel $3.6309$ $3.7486$ $3.7686$
Wiener filter $3.8878$ $4.0373$ $4.2206$
Adaptive CCTFD $\mathbf{4.1225}$ $\mathbf{4.2956}$ $\mathbf{4.4592}$
@@ 5 Conclusion
The convolution type of CCTFD time-frequency analysis method suitable for the adaptive filter denoising for additive noises jamming signals under the condition of low SNR was established, giving rise to the adaptive kernel function which minimizes the MSE in the WVD domain. The proposed adaptive CCTFD can automatically adjust its kernel function according to the change of signal to adapt to different signal characteristics. It turns out that its denoising effect is superior not only to some classical fixed kernel function-based CCTFDs, including the Margenau-Hill distribution, the Kirkwood-Rihaczek distribution, the Born-Jordan distribution, and the Page distribution, but also to the adaptive radial Gaussian kernel function-based CCTFD. Under white Gaussian noise, our method outperforms the ordinary Wiener filter, while under colored noise, it performs comparably to the Wiener filter. Experiments on real-world radar data from the RadChar dataset further confirm that the proposed method achieves superior MSE and PSNR, demonstrating its practical applicability and robustness.
@@
[…]
Author contributions M.J.C. conceived the study, developed the methodology, performed the experiments, and drafted the manuscript. Z.Z.C. supervised the research, contributed to the theoretical analysis, and provided critical revisions to the manuscript. All authors reviewed and approved the final version of the manuscript.
[…]
Funding This research is supported by the Open Foundation of Hubei Key Laboratory of Applied Mathematics (Hubei University) (HBAM202404), and the Key Laboratory of System Control and Information Processing, Ministry of Education (Scip20240121).
[…]
Data availability The experiment data used to support the findings of this study are available from the corresponding author upon request.
Declarations
[…]
Competing interests The authors declare no competing interests.